# Heterogeneous Intertemporal Treatment Effects via Dynamic Panel Data Models

### v1

@@ 3 Causal Inference via Dynamic Panel Data Methods
3.1 Dynamic Panel Data Models
Canonical dynamic panel data methods are rooted in linear models with lagged outcomes as a regressor. In our context, we consider causal interpretations of a dynamic panel data model (DPDM) for observed outcomes:
@@ 3.1 Dynamic Panel Data Models
where $\gamma,\beta$ are constant parameters, $\alpha_{i}$ are time-invariant individual fixed effects, $\theta_{t}$ are deterministic time trends, and $\varepsilon_{it}$ are time-varying idiosyncratic errors. Specifically, we investigate how the estimated coefficients can be related to the moments of heterogeneous treatment effects and trends in potential outcomes.
@@
Equation (7) is natural for empiricists as a reduced-form model for a series of observed outcomes. For example, a natural interpretation is an extension of a static panel data model in which the researcher also wishes to control for lagged outcomes. A challenge for estimating (7) is the endogeneity of lagged outcomes on the right-hand side, due to the unobservable individual fixed effects $\alpha_{i}$. The ordinary least-squares (OLS) estimators and fixed-effect estimators using either “within” transformation or first-differences are biased and inconsistent when applied to (7) (e.g., (Baltagi, 2021)). Anderson and Hsiao (1982) and Arellano and Bond (1991) proposed consistent estimators for the coefficients in (7), using moments implied by instruments such as the lagged outcomes and covariates.
Next, we turn to the interpretation of such estimators. It is important to emphasize, however, that we do not interpret the DPDM for observed outcomes (7) as a causal model per se. Rather, we consider its implicit interpretation and estimation under non-parametric assumptions on the potential outcome process. This exercise is analogous to the interpretation of TWFE regressions under parallel trend assumptions, e.g., De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021). In Section 3.4 we also consider an analogous parametric model to (7), but as a model of the potential outcome generating process.
3.2 Instrumented First-Difference Estimators
Our goal in this section is to obtain a causal interpretation of estimators for (7) that address the endogeneity of lagged outcomes. First-difference transformation of (7) implies:
$\Delta Y_{it}=\gamma\Delta Y_{it-1}+\beta\Delta D_{it}+\Delta\theta_{t}+\Delta\varepsilon_{it}\ \quad\text{for }2\leq t\leq T\text{,}$ (8)
@@ 3.2 Instrumented First-Difference Estimators
where $\Delta$ denotes the first-difference operator, e.g., $\Delta Y_{it}\equiv Y_{it}-Y_{it-1}$. Anderson and Hsiao (1982) estimated the coefficients via IV regression, using $Y_{it-2}$ as instruments for $\Delta Y_{it-1}$ when $D_{it}$ is strictly exogenous. Arellano and Bond (1991) proposed a GMM estimator when $D_{it}$ is sequentially exogenous (a.k.a. pre-determined), using earlier lagged outcomes and treatments as instruments for $\Delta Y_{it-1}$ and $\Delta D_{it}$.
We focus henceforth on the causal interpretation of a two-stage least squares (2SLS) estimator for $\beta$ when $T=2$. In this case, the coefficients $\rho$ and $\gamma$ are just identified, with $(Y_{i0},D_{i1})$ being instruments for $(\Delta Y_{i1},\Delta D_{i2})$ in (8), and 2SLS and GMM estimators are numerically equivalent. As before, we drop individual subscripts $i$ for simplicity.
@@
$\tau_{1}\equiv Y_{1}(1)-Y_{1}(0);\quad\tau_{2}(d_{1})\equiv Y_{2}(d_{1},1)-Y_{2}(d_{1},0).$ (10)
The trend of observed outcomes can then be written as:
@@
To fix ideas, suppose $Y_{0}$ takes discrete values. Let the first-stage regression be fully saturated, so that the first-stage fitted values of the trend in observed outcomes and treatments converge in probability to $E[\Delta Y_{1}|Y_{0},D_{1}]$ and $E[\Delta D_{2}|Y_{0},D_{1}]$ respectively.
The next lemma invokes the Frisch-Waugh-Lovell Theorem and the decomposition of $\Delta Y_{2}$ in (12) to express the probability limit of the 2SLS estimator for $\beta$ in terms of causal and non-causal components. Let $w(Y_{0},D_{1})$ be the projection errors in the linear projection of $E(\Delta D_{2}|Y_{0},D_{1})$ on $E(\Delta Y_{1}|Y_{0},D_{1})$. Assume $w(Y_{0},D_{1})$ is not degenerate at zero, which is an innocuous, empirically verifiable condition.
@@
Suppose Assumption 1 holds and $w(Y_{0},D_{1})$ is not degenerate at zero. Let the first-stage regressions be fully saturated. The 2SLS estimator of $\beta$ in (8), which uses $(Y_{0},D_{1})$ as instruments for $(\Delta Y_{1},\Delta D_{2})$, converges in probability to:
@@
Lemma 1 clarifies the basis and impediments to causal interpretations of the 2SLS estimator with no other conditions than the parametrization of potential outcomes (Assumption 1). It does not invoke sequential exchangeability in Assumption 2. The first term on the right in (13) is a weighted aggregate of treatment effects in period 2 among the treated ($D_{2}=1$), yet with no guarantee that the combination is convex or even positively weighted.333 This is analogous to a point by De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021) in the causal interpretation of two-way fixed-effect (TWFE) regressions in static panel data models. The second term, if non-zero, is an additional confounder to causal interpretation in terms of treatment effects $\tau_{2}(d_{1})$.
To endow the 2SLS estimand in (13) with a clearer causal interpretation, we require further restrictions on heterogeneous treatment effects in addition to those on selection in Assumption 2.
@@
Note first that each of the conditions in Assumption 3 would be satisfied in the special case where treatment effects $\tau_{t}(\cdot)$ and untreated trends $\delta_{t}(\cdot)$ are homogeneous in the population. Under Assumption 3, the next proposition shows that the 2SLS first-difference estimator recovers a convex combination of history-dependent treatment effects.
@@
Suppose Assumptions 1, 2a, 3 hold, and $w(Y_{0},D_{1})$ is not degenerate at zero. The 2SLS estimator of $\beta$ in (8) converges in probability to a convex combination of treatment effects:
@@
This proposition confirms that an intuitive causal interpretation of the 2SLS estimand for the coefficient of $\Delta D_{2}$ exists under sufficiently restrictive conditions. A proof is presented in Appendix A. Intuitively, the proof proceeds in the three steps. First, derive the expression in (19) for the projection errors. Next, eliminate the confounding (second) term in (13). Finally, establish the convex weights in (18).
A technical note worth mentioning is that only the mean version of sequential exchangeability (Assumption 2b) is invoked in the first and second steps of the proof; in the third step, the full version (Assumption 2a) is required in period $t=1$ to separate restrictions on selection from restrictions on heterogeneity.
Proposition 2 complements the existing literature. It is well-known that the average treatment effect $E[\tau(d_{1})]$ is identified for $d_{1}\in\{0,1\}$ under sequential exchangeability in Assumption (2); see, e.g., an application of the $g$-formula in (Robins and Hernán, 2008). In comparison, our contribution is to show how to use an alternative 2SLS approach to estimate a different causal effect parameter, which is a weighted aggregate of treatment effects where the weights can be recovered from observable moments.
3.3 An Adjusted Inverse-Propensity-Weighted Estimator
The result in Proposition 2 naturally raises a new question: can we modify our approach to construct new estimators for other averages of causal effects, without restricting the heterogeneity of treatment effects as in Assumption 3?
Our next result shows how to use an adjusted inverse-propensity-weighted (IPW) method to estimate an average treatment effect (ATE) upon transforming $\Delta Y_{2}$. The basis for this transformation is the same decomposition (12) invoked in the preceding subsection 3.2. In our dynamic potential outcome setting, treatment effects are also indexed by past treatment $\tau_{2}(d_{1})$. Thus, the ATE in period 2 that we target also averages over the observed past treatment $D_{1}$. This new (to the best of our knowledge) estimator does not invoke the restrictions on heterogeneous treatment effects and trends in Assumption 3 or the full version of sequential exchangeability of Assumption 2a; besides Assumptions 1 and 2b, it only requires a standard support assumption on treatment paths.
@@ 3.3 An Adjusted Inverse-Propensity-Weighted Estimator
$E[D_{2}|Y^{1},D_{1}]\in(0,1)$.
Under Assumption 4, define:
@@
The next proposition establishes that the mean of $\widetilde{\Delta Y_{2}}$ identifies the ATE in period 2, given the realized treatments in period 1.

### v2

@@ 3 Causal Inference via Dynamic Panel Data Methods
3.1 DPDM for Observed Outcomes
We consider causal interpretations of an estimand motivated by an instrument-based method for estimating a dynamic panel data model (DPDM) for observed outcomes:
@@ 3.1 DPDM for Observed Outcomes
where $\gamma,\beta$ are constant parameters, $\alpha_{i}$ are time-invariant individual fixed effects, $\theta_{t}$ are deterministic time trends, and $\varepsilon_{it}$ are time-varying idiosyncratic errors. We investigate how the estimated coefficients can be related to the moments of heterogeneous treatment effects and trends in potential outcomes.
@@
Equation (3) is natural for empiricists as a reduced-form model for a series of observed outcomes. For example, a natural interpretation is an extension of a static panel data model in which the researcher also wishes to control for lagged outcomes. A challenge for estimating (3) is the endogeneity of lagged outcomes on the right-hand side, due to the unobservable individual fixed effects $\alpha_{i}$. The ordinary least-squares (OLS) estimators and fixed-effect estimators using either “within” transformation or first-differences are biased and inconsistent when applied to (3) (e.g., (Baltagi, 2021)).
Anderson and Hsiao (1982) and Arellano and Bond (1991) proposed consistent estimators for the coefficients in (3), using moments implied by instruments such as the lagged outcomes and covariates. Specifically, first-differencing (3) implies:
$\Delta Y_{it}=\beta\Delta D_{it}+\gamma\Delta Y_{it-1}+\Delta\theta_{t}+\Delta\varepsilon_{it}\ \quad\text{for }2\leq t\leq T\text{,}$ (4)
@@
where $\Delta$ denotes the first-difference operator, e.g., $\Delta Y_{it}\equiv Y_{it}-Y_{it-1}$. Anderson and Hsiao (1982) estimated the coefficients via IV regression, using $Y_{it-2}$ as instruments for $\Delta Y_{it-1}$ when $D_{it}$ is strictly exogenous. Arellano and Bond (1991) proposed a GMM estimator when $D_{it}$ is sequentially exogenous (a.k.a., pre-determined), using earlier lagged outcomes and treatments as instruments for $\Delta Y_{it-1}$ and $\Delta D_{it}$.
It is important to emphasize that we do not interpret the DPDM for observed outcomes (3) as a causal model per se. Rather, our goal is to obtain a causal interpretation of the limits to which a first-differenced instrument-based estimator converge. We refer to these limits as the Arellano-Bond (first-differenced) instrumental-variable (IV) estimands of DPDMs (or DPDEs). We show how these estimands relate to the distribution of heterogeneous treatment effects under non-parametric assumptions on the potential outcome process. This exercise is analogous to the interpretation of TWFE regressions under parallel trend assumptions, e.g., De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021).
3.2 Causal Interpretation of First-Differenced IV Estimands
We focus henceforth on the causal interpretation of IV estimands of DPDM in (3) when $T=2$, which are defined as follows (we drop individual subscripts $i$ for simplicity):
$\begin{pmatrix}\widetilde{\beta}\\ \widetilde{\gamma}\\ \widetilde{\Delta\theta_{2}}\end{pmatrix}\equiv E[Z^{\prime}X]^{-1}E[Z^{\prime}\Delta Y_{2}],$ (5)
where
$X$ $\equiv(\Delta D_{2},\Delta Y_{1},1)$
$Z$ $\equiv E[X|Y_{0},D_{1}]$
are each row vectors.
If equation (3) is the actual data-generating process (DGP) for the observed outcomes, $\widetilde{\beta}$ is the probability limit of an IV estimator for $\beta$ with $(E[\Delta D_{2}|Y_{0},D_{1}],E[\Delta Y_{1}|Y_{0},D_{1}])$ instrumenting for $(\Delta D_{2},\Delta Y_{1})$.555 This IV estimator would also be numerically equivalent to a 2SLS estimator that uses $(D_{1},Y_{0})$ as instruments for $(\Delta D_{2},\Delta Y_{1})$, if the conditional expectations of $\Delta D_{2},\Delta Y_{1}$ are linear in $(Y_{0},D_{1})$. With such a choice of instruments, the model is just-identified, with GMM/2SLS estimators numerically identical to the IV estimator. In addition, if the DGP further satisfies the identifying assumptions in Arellano and Bond (1991), such as exogeneity of $D_{it}$ and serial uncorrelation of reduced-form errors $\varepsilon_{it}$, then $\widetilde{\beta}$ would indeed coincide with $\beta$ in (3).
A natural question to ask is the following: are there any DGPs of potential outcomes and treatments that would lead to a model for observed outcomes that conform to the functional form in (3) and satisfy the identifying conditions in Arellano and Bond (1991) (so that $\widetilde{\beta}=\beta$)? The answer is positive; we present it in Section 3.3.1.
For the rest of Section 3.2, we address a different question, namely, whether the IV estimand $\widetilde{\beta}$ above admits a causal interpretation in general. We explore conditions under which it does. While doing so, we maintain that the DGP of observed outcomes does not necessarily conform to (3) or the identifying conditions in Arellano and Bond (1991).
@@ 3.2 Causal Interpretation of First-Differenced IV Estimands
$\tau_{1}\equiv Y_{1}(1)-Y_{1}(0);\quad\tau_{2}(d_{1})\equiv Y_{2}(d_{1},1)-Y_{2}(d_{1},0).$
The trend of observed outcomes are decomposed as:
@@
The next lemma invokes the Frisch-Waugh-Lovell Theorem and the decomposition of $\Delta Y_{2}$ in (8) to express the IV estimand $\widetilde{\beta}$ in (5) in terms of causal and non-causal components. Let $w(Y_{0},D_{1})$ be the errors in the linear projection of $Z_{1}\equiv E(\Delta D_{2}|Y_{0},D_{1})$ onto $Z_{-1}\equiv(E(\Delta Y_{1}|Y_{0},D_{1}),1)$:
$w(Y_{0},D_{1})\equiv Z_{1}-Z_{-1}E[Z_{-1}^{\prime}Z_{-1}]^{-1}E[Z_{-1}^{\prime}Z_{1}].$ (9)
Assume $w(Y_{0},D_{1})$ is not degenerate at zero, which is an innocuous, empirically verifiable condition.
@@
Suppose Assumption 1 holds, and $w(Y_{0},D_{1})$ as defined in (9) is not degenerate at zero. The IV estimand $\widetilde{\beta}$ in (5) is:
@@
Lemma 1 clarifies the basis and impediments to causal interpretations of the IV estimand $\widetilde{\beta}$ with no other conditions than Assumption 1. It does not invoke sequential exchangeability in Assumption 2. The first term on the right in (10) is a weighted aggregate of individually heterogeneous treatment effects $\tau_{2}(\cdot)$ in the second period among the treated ($D_{2}=1$), yet with no guarantee that the combination is convex or even positively weighted.666 This is analogous to a point by De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021) in the causal interpretation of two-way fixed-effect (TWFE) regressions in static panel data models. The second term, if non-zero, is an additional confounder to causal interpretation in terms of treatment effects $\tau_{2}(\cdot)$.
To endow this decomposition of the IV estimand $\widetilde{\beta}$ in (10) with a clearer causal interpretation, we require further restrictions on heterogeneous treatment effects in addition to those on selection in Assumption 2.
@@
Assumption 3 requires the first moments of treatment effects $\tau_{t}(\cdot)$ and untreated trends $\delta_{t}(\cdot)$ to be invariant in the initial condition or earlier potential outcomes; obviously, it holds when these treatment effects and untreated trends are homogeneous in the population.
The next proposition shows that under Assumption 3 the IV first-difference estimand recovers a convex combination of history-dependent treatment effects.
@@
Suppose Assumptions 1, 2a, 3 hold, and $w(Y_{0},D_{1})$ is not degenerate at zero. Then the IV estimand $\widetilde{\beta}$ in (5) is a convex combination of treatment effects:
@@
This proposition confirms that an intuitive causal interpretation of the IV estimand for the coefficient of $\Delta D_{2}$ exists under additional conditions on individual heterogeneity. Intuitively, the proof (presented formally in Appendix A) proceeds in three steps. First, derive the expression of the projection errors in (16). Next, eliminate the confounding (second) term in (10). Finally, establish the convex weights in (15).
Only the mean version of sequential exchangeability (Assumption 2b) is invoked in the first and second steps; in the third step, the full version (Assumption 2a) is required in period $t=1$ to separate restrictions on selection from restrictions on heterogeneity.
Proposition 2 complements the existing literature. It is well-known that the average treatment effect (ATE) $E[\tau_{2}(d_{1})]$ is identified for $d_{1}\in\{0,1\}$ under sequential exchangeability in Assumption (2); see, e.g., the $g$-formula in (Robins and Hernán, 2008).
In comparison, in Proposition 2 we propose an alternative way to use the sequential exchangeability condition. By focusing only on causal interpretation of first-differenced IV estimators (instead of a more ambitious goal of estimating $E[\tau_{2}(d_{1})]$ for all $d_{1}$), we manage to obtain useful insights under additional mean restrictions on the heterogeneous treatment effects (Assumption 3).
We conclude the section by noting that an outcome-modifying mean estimand identifies the average treatment effect $E[\tau_{2}(D_{1})]$. This is a different causal parameter than the one targeted in the $g$-formula. This result does not require the restrictions on heterogeneous treatment effects and trends in Assumption 3 or the full sequential exchangeability of Assumption 2a; besides Assumptions 1 and 2b, it only requires a standard support assumption on treatment paths below.
@@
$E[D_{2}|Y^{1},D_{1}]\in(0,1)$ almost surely.
Under Assumption 4, define:777 Note that $\widetilde{\Delta Y_{2}}$ is identical to an expression replacing terms $\Delta Y_{2}$ with $Y_{2}$. We preserve this redundant trend form in order to clarify the connection to our first-differenced IV estimand.
@@
Relative to the decomposition (8), the transformation from $\Delta Y_{2}$ to $\widetilde{\Delta Y_{2}}$ in expectation removes the confounder $\delta_{2}(D_{1})$ through the term $E[\Delta Y_{2}|Y^{1},D_{1},D_{2}=0]$ and weights by the treated observations $E[D_{2}|Y^{1},D_{1}]$. The next proposition establishes that the mean of $\widetilde{\Delta Y_{2}}$ identifies the ATE in period 2, given the realized treatments in period 1.